# Periodic orbits and their gravitational waves in EMRIs: supermassive black hole affected by galactic dark matter halos

### v2

@@ II.3 Periodic Orbits and Their Precession Parameter $q$
III Effects of Dark Matter Halos on Periodic Orbits
In this section, we present numerical results on periodic orbits, focusing on the dark matter halo effects on the shape of these orbits. For a given periodic orbit characterized by a group of zoom, whirl, vertex configurations (zwv)(z~w~v), its shape and trajectory are determined by the conserved energy $E$ and orbital angular momentum $L$ of a massive particle. A proper choice of $(E,L)$ in their permissible ranges, which are prescribed by Eq. (16), could result in a periodic orbit around the SMBH with integers (zwv)(z~w~v).To systematically investigate the effects of dark matter halos on the precession angle and periodic orbits’ shape, we can either maintain the orbital angular momentum $L$ and change the energy (which leads to a series $E_{(z~w~v)}$), or maintain the energy as constant and vary the angular momentum (which results in a series $L_{(z~w~v)}$). In this section, we maintain the angular momentum to give comparison of periodic orbits calculated from different dark matter halo models, while relegating the case of selected energy $E$ to Appendix A.Specifically, we let the angular momentum $L$ satisfy the relation
$L=L_{\text{ISCO}}+\varepsilon(L_{\text{MBO}}-L_{\text{ISCO}}),$ (24)
where $\varepsilon\in[0,1]$, such that the permissible range in Eq. (16) is automatically satisfied. For the sake of simplicity, we choose $\varepsilon=0.5$ for angular momentum in this section, while relegating the case of selected energy $E$ to Appendix A.
Furthermore, as demonstrated by Eqs. (6a)–(6c), the spacetime metric of the gravitational field is determined by the dark matter mass $k$ and the halo characteristic radius parameter $h$. The geometric shapes of periodic orbits around SMBH could be inevitably affected by these parameters. Therefore, in the remaining part of this work, we primarily examine two scenarios: A. the effect of dark matter mass; B. the impact of dark matter halo scale.
III.1 The Effect of Dark Matter Mass on Periodic Orbits
In the study of dark matter effects on periodic orbits (which is conducted under a selected angular momentum $L=L_{\text{ISCO}}+\frac{1}{2}(L_{\text{MBO}}-L_{\text{ISCO}})$), the primary step is to determine the characteristic energy $E_{(z~w~v)}$.The accurate value of this energy can be calculated from the numerical results of the precession angle parameter $q$ varying with $E$, as is derived from Eq. (21). Fig. 4 exhibits the variation patterns of precession angles under three different dark matter halo models (NFW, Beta, and Moore) with different dark matter masses $k$, where the dark matter halo scale is fixed at $h=10^{7}M$. Fig. 5 further compares the precession angle among different models under fixed dark matter mass, and halo scale. The energy $E_{(z~w~v)}$ corresponding to periodic orbits can be determined via the intersection points of the curves with $q=w+v/z$, and the obtained precise values of $E_{(z~w~v)}$ is detailed in Appendix B.
@@ III.1 The Effect of Dark Matter Mass on Periodic Orbits
Figure 5: Comparison of precession angles for different dark matter halo models: the dark matter halo scale is fixed at $h=10^{7}M$, and the subfigures illustrate results under four distinct dark matter masses, including (a) $k=1\times 10^{3}M$; (b) $k=3\times 10^{3}M$; (c) $k=1\times 10^{4}M$; and (d) $k=2\times 10^{4}M$.
@@
After obtaining the corresponding orbital energy parameters $E_{(z~w~v)}$ under different dark matter masses $k$, we can solve for the trajectories of periodic orbits and systematically investigate the influence of dark matter on the geometric shapes of these periodic orbits. Fig. 6 shows the periodic orbits around a Schwarzschild black hole and periodic orbits around black holes embedded in three different dark matter halos. The layout of Fig. 6 is as follows: each row corresponds to the same set of orbital configurations (zwv)(z~w~v), presenting the periodic orbits near the Schwarzschild black hole and those in three dark matter halo environments; each column fixes the dark matter mass $k$, exhibiting the variation pattern of orbital shapes by varying the orbital configurations (zwv)(z~w~v). From horizontal comparative analysis, it is shown that the presence of dark matter halos significantly stretches the apoapsis of periodic orbits. As the dark matter mass $k$ increases, the deviation between orbits in dark matter halo environments and Schwarzschild case gradually intensifies. The physical reason behind this phenomenon stems from the fact that, for the same precession angle $q$, the larger dark matter mass, the differences between orbital energies (dark matter halo environments vs Schwarzschild case) are more pronounced (as shown in Fig. 4). Notably, the NFW and Beta models exhibit nearly identical orbital trajectories for all dark matter masses $k$ considered in this study, which is consistent with their overlapping in precession angle curves shown in Fig. 5. This demonstrates that despite the different functional forms of NFW and Beta density profiles, they produce nearly indistinguishable effects on both the energetics and shapes of periodic orbits. Furthermore, Fig. 6 demonstrates that periodic orbits calculated in NFW and Beta models are relatively more extended than those in Moore model and Schwarzschild black hole cases. This is caused by the fact that the orbital energies calculated in the NFW and Beta models are substantially lower than those in the Moore model for the same precession angle q (as illustrated in Figs. 4 and 5), since orbital energy is inversely related to the apoapsis distance. This finding suggests a stronger gravitational influence from the NFW/Beta dark matter halos, which is consistent with the results for the ISCO and MBO in II.2.
@@
Furthermore, to isolate the influences of dark matter halo models, we construct periodic orbits within the NFW model using different values of $k$, each combined with five distinct orbital configurations (zwv)(z~w~v). The results, presented in Fig. 7, clearly demonstrate the effect of dark matter mass k on periodic orbital shapes. For example, the relationship between apoapsis distance and dark matter mass exhibits a consistent monotonic trend: an increase in dark matter mass leads to further apoapsis distance and more pronounced differences of orbital shape, compared with Schwarzschild case. Additionally, from the comparative analysis in Figs. 6 and 7, we observe that the orbital configurations (zwv)(z~w~v) directly determine the geometric shape of periodic orbits. Specifically, (zwv)(z~w~v) represent the number of “leaves”, the count of additional whirls, and the behavior of the subsequent vertex reached by the particle after leaving the initial vertex (apoapsis). As a result, larger parameter values correspond to more complex orbital structures, which is consistent with the zoom-whirl-vertex classification of periodic orbits in pioneering studies [38, 39].
III.2 The Effect of Dark Matter Halo Scale on Periodic Orbits
@@
IV Gravitational waveforms from periodic orbits
Extreme Mass Ratio Inspiral (EMRI) systems are key target sources for future space-based gravitational wave detectors. These systems are formed by stellar-mass compact objects orbiting around supermassive black holes, and their gravitational wave signals contain rich information about the system’s dynamics and the spacetime geometry of the central black hole. When a small object orbits around a supermassive black hole enveloped by dark matter (DM) in a periodic motion, the gravitational waves it emits provide a unique avenue for studying the system’s properties. This section provides an analysis of the gravitational waveforms of EMRI systems generated by periodic motions, and the theoretical framework for calculating gravitational waves from such periodic orbits is also briefly reviewed. In the calculation of gravitational waves in EMRI systems, the adiabatic approximation and “Numerical Kludge” waveform model, which is applicable when the small object moves in a nearly static gravitational background with circular, precessional or periodic motion, is adopted [111, 112, 113, 114, 115, 116, 117, 118, 119, 120]. Under this approximation, the energy and angular momentum of the smaller object can be considered constant over several orbital periods, and its trajectory can be viewed as a geodesic in the static background spacetime, with the gravitational radiation’s reaction on the object’s motion temporarily ignored.
The “Numerical Kludge” waveform model provides a practical scheme for calculating gravitational waves from periodic orbits in a DM halo environment [121]. This method consists of two steps: first, the motion equations Eqs. 10a , 10b and 12 are numerically solved to determine the orbit of the small body in the gravitational spacetime containing the DM distribution; then, the quadrupole formula for gravitational radiation is applied to this orbit to generate the corresponding waveform [122]. This method can conveniently reveal the gravitational wave signals of a slowly evolving EMRI system, providing possibilities for exploring the characteristics of the periodic orbits (or precessional orbits), central black hole, and surrounding DM distribution. For the spacetime metric perturbation $h_{ij}$, the gravitational radiation quadrupole moment formula calculated to second order can be expressed as [123, 124, 48],
$h_{ij}=\frac{4\mu M}{D_{L}}\left(v_{i}v_{j}-\frac{m}{r}n_{i}n_{j}\right),$ (25)
@@ IV Gravitational waveforms from periodic orbits
$\mathbf{e}_{X}$ $=(\cos\zeta,-\sin\zeta,0),$ (26a)
$\mathbf{e}_{Y}$ $=(\sin\iota\sin\zeta,\cos\iota\cos\zeta,-\sin\iota),$ (26b)
$\mathbf{e}_{Z}$ $=(\sin\iota\sin\zeta,-\sin\iota\cos\zeta,\cos\iota).$ (26c)
In general relativity, the polarization state of gravitational waves is typically decomposed into two independent modes: “+” (orthogonal) polarization and “×” (cross) polarization. In the introduced coordinate system, these two polarization components $h_{+}$ and $h_{\times}$ detected by observers can be expressed as [48, 65],
$h_{+}$ $=\frac{1}{2}(e^{i}_{X}e^{j}_{X}-e^{i}_{Y}e^{j}_{Y})\times h_{ij}=-\frac{2\mu M^{2}}{D_{L}r}\left(1+\cos^{2}\iota\right)\cos\left(2\phi+2\zeta\right)$ (27a)
$h_{\times}$ $=\frac{1}{2}(e^{i}_{X}e^{j}_{Y}+e^{i}_{Y}e^{j}_{X})\times h_{ij}=-\frac{4\mu M^{2}}{D_{L}r}\cos\iota\sin\left(2\phi+2\zeta\right),$ (27b)
@@
IV.1 The effect of dark matter mass on gravitational waves
@@
IV.2 The effect of dark matter halo scale on gravitational waves
@@ V Conclusion
This study systematically investigates the influence of dark matter halos on particles’ periodic orbits around a black hole (with zoom-whirl-vertex behavior) and the produced gravitational wave radiation. This work is carried out within the framework of spherically symmetric dark matter halo models, which have been widely used in astrophysical studies. Through detailed analysis of three dark matter halo models—NFW, Beta, and Moore—we examine how dark matter halo parameters affect spacetime metrics, effective potentials, and properties of periodic orbits around black holes, as well as the resulting gravitational wave signals.
@@
The effects of dark matter halos on spacetime geometry, particle orbits and gravitational waves can be incorporated into two parameters: the dark matter mass $k$ and halo characteristic radius $h$. Our analysis of the effective potential reveals that the existence of dark matter halos reduces the extrema of the effective potential. This reduction becomes more obvious as the dark matter mass $k$ increases, directly affecting the MBO and ISCO of particles. The presence of dark matter halos increases the angular momentum for both MBO and ISCO, with the NFW and Beta models showing nearly identical behavior, while the Moore model exhibits lower angular momentum values.
@@
To comprehensively explore periodic orbits, we examine the effects of both dark matter mass $k$ and halo characteristic radius $h$ under the condition of selected angular momentum $L=L_{\text{ISCO}}+\varepsilon(L_{\text{MBO}}-L_{\text{ISCO}})$ and the selected energy $E$ (in Appendix B). For the selected angular momentum condition, it is demonstrated that for a given precession angle $q$, larger dark matter masses require lower orbital energies $E$. The NFW and Beta models require substantially lower energies $E_{(z~w~v)}$ than the Moore model for the same precession angle and orbital configurations (zwv)(z~w~v). This energy difference directly affects the orbital shapes, increasing dark matter mass stretches the apoapsis of periodic orbits, which leads to larger orbital dimensions. For a fixed precession angle $q$, larger dark matter halo scales require higher orbital energy. At smaller scales, periodic orbital shapes differ from the Schwarzschild case. As the halo scale increases, the orbits gradually converge toward the Schwarzschild trajectories, closely approaching them at large scales. This behavior reflects the dilution of gravitational effects: when the dark matter mass remains constant while the halo scale increases, the gravitational binding weakens in the regions. Particles must then possess higher energies (closer to the asymptotic value $\lim_{r\to\infty}V_{\text{eff}}=1$ without gravitation) to maintain the same precession characteristics. This scale-dependent effect is clearly shown and explained with five different periodic orbit configurations (zwv)(z~w~v), where orbits gradually approach the Schwarzschild case as $h$ increases. In particular, when the dark matter halo scale is $h\sim 10^{10}M$, the periodic orbits of all three dark matter models become indistinguishable from the Schwarzschild reference, exhibiting the dilution limit of dark matter halo effects. Furthermore, analysis of periodic orbits with fixed $E$ yields conclusions consistent with those obtained for fixed $L$.
@@
To explore the observational signatures of these periodic orbits, we analyze an extreme-mass-ratio inspiral (EMRI) system consisting of a test object with a typical mass ratio and luminosity distance. The gravitational wave analysis shows strong correspondence between the waveforms and orbital properties. Larger dark matter mass leads to more pronounced differences in generated gravitational waves between periodic orbits around SMBH in dark matter halos and those of Schwarzschild black holes, enhancing the distinguishability among three dark matter halo models. The waveforms exhibit stretching trends, particularly an increase in the duration of each complete orbital period, which stems directly from the modification of the gravitational potential by the dark matter halo. In contrast to the dark matter mass effects, a larger dark matter halo scale $h$ results in a more diffuse mass distribution, weakening its effect on the gravitational potential. The gravitational waveforms generated by periodic orbits exhibit significant differences between various dark matter halo models under a relatively small halo scale, which can be used to constrain dark matter models with future observations. The gravitational waveforms gradually approach those for Schwarzschild black holes when dark matter halo scales become sufficiently large. Particularly, when the characteristic radius of the dark matter halo is $h\sim 10^{10}M$, the periodic orbits and their gravitational waveforms converge precisely with those of the Schwarzschild black hole. Furthermore, examining the waveform structure reveals that the orbital configurations (zwv)(z~w~v), as fundamental quantities characterizing the geometric shapes of periodic motion, directly determine the spatiotemporal features (the zoom-whirl stages) of the gravitational wave signals.
@@
Throughout our analysis, we observe that the NFW and Beta models produce nearly indistinguishable results for both orbital characteristics and gravitational wave signals for all parameter ranges examined in this study. This accordance holds for precession angles, periodic orbit configurations, and gravitational waveforms, suggesting that these two density profiles have very similar effects on particle orbits and gravitational radiation in the parameter ranges relevant to galactic dark matter halos. The Moore model, while following similar qualitative trends, generates different quantitative results compared with Beta and NFW models, which become more obvious at larger dark matter masses.
@@
In summary, through an systematic analysis of dark matter halo models, this study clarifies how dark matter halo parameters influence particle orbits around black holes and gravitational wave radiation. The complementary effects of the dark matter mass and its characteristic radius offer a pathway to understand the complex interplay between dark matter halo and SMBH in galactic center, as well as provide theoretical predictions for gravitational waveforms that can be compared with observations. These findings may provide useful constraints for interpreting gravitational wave signals from EMRIs detected by future space-based observatories, such as LISA, TianQin, and Taiji.
Acknowledgements.

### v3

@@ 2.3 Periodic Orbits and Their Precession Parameter $q$
3 Effects of Dark Matter Halos on Periodic Orbits
In this section, we present numerical results on periodic orbits, focusing on the dark matter halo effects on the shape of these orbits. For a given periodic orbit characterized by a group of zoom, whirl, vertex configurations (zwv)(z~w~v), its shape and trajectory are determined by the conserved energy $E$ and orbital angular momentum $L$ of a massive particle. A proper choice of $(E,L)$ in their permissible ranges, which are prescribed by Eq. (2.16), could result in a periodic orbit around the SMBH with integersă(zwv)(z~w~v). To systematically investigate the effects of dark matter halos on the precession angle and periodic orbits’ shape, we can either maintain the orbital angular momentum $L$ and change the energy (which leads to a series $E_{(z~w~v)}$), or maintain the energy as constant and vary the angular momentum (which results in a series $L_{(z~w~v)}$). In this section, we maintain the angular momentum to give comparison of periodic orbits calculated from different dark matter halo models, while relegating the case of selected energy $E$ to Appendix A. Specifically, we let the angular momentum $L$ satisfy the relation
$L=L_{\text{ISCO}}+\varepsilon(L_{\text{MBO}}-L_{\text{ISCO}}),$ (3.1)
where $\varepsilon\in[0,1]$, such that the permissible range in Eq. (2.16) is automatically satisfied. For the sake of simplicity, we choose $\varepsilon=0.5$ for angular momentum in this section, while relegating the case of selected energyăEEăto Appendix A.
Furthermore, as demonstrated by Eqs. (2.6a)-(2.6c), the spacetime metric of the gravitational field is determined by the dark matter mass $k$ and the halo characteristic radius parameter $h$. The geometric shapes of periodic orbits around SMBH could be inevitably affected by these parameters. Therefore, in the remaining part of this work, we primarily examine two scenarios: A. the effect of dark matter mass; B. the impact of dark matter haloăscale.
3.1 The Effect of Dark Matter Mass on Periodic Orbits
In the study of dark matter effects on periodic orbits (which is conducted under a selected angular momentum $L=L_{\text{ISCO}}+\frac{1}{2}(L_{\text{MBO}}-L_{\text{ISCO}})$), the primary step is to determine the characteristic energy $E_{(z~w~v)}$.The accurate value of this energy can be calculated from the numerical results of the precession angle parameter $q$ varying with $E$, as is derived from Eq. (2.21). Fig. 4 exhibits the variation patterns of precession angles under three different dark matter halo models (NFW, Beta, and Moore) with different dark matter masses $k$, where the dark matter halo scale is fixed at $h=10^{7}M$. Fig. 5 further compares the precession angle among different models under fixed dark matter mass, and halo scale. The energy $E_{(z~w~v)}$ corresponding to periodic orbits can be determined via the intersection points of the curves with $q=w+v/z$, and the obtained precise values of $E_{(z~w~v)}$ is detailed in Appendix B.
@@ 3.1 The Effect of Dark Matter Mass on Periodic Orbits
[…]
@@
After obtaining the corresponding orbital energy parameters $E_{(z~w~v)}$ under different dark matter masses $k$, we can solve for the trajectories of periodic orbits and systematically investigate the influence of dark matter on the geometric shapes of these periodic orbits. Fig. 6 shows the periodic orbits around a Schwarzschild black hole and periodic orbits around black holes embedded in three different dark matter halos. The layout of Fig. 6 is as follows: each row corresponds to the same set of orbital configurations (zwv)(z~w~v), presenting the periodic orbits near the Schwarzschild black hole and those in three dark matter halo environments; each column fixes the dark matter mass $k$, exhibiting the variation pattern of orbital shapes by varying the orbital configurations (zwv)(z~w~v). From horizontal comparative analysis, it is shown that the presence of dark matter halos significantly stretches the apoapsis of periodic orbits. As the dark matter mass $k$ increases, the deviation between orbits in dark matter halo environments and Schwarzschild case gradually intensifies. The physical reason behind this phenomenon stems from the fact that, for the same precession angle $q$, the larger dark matter mass, the differences between orbital energies (dark matter halo environments vs Schwarzschild case) are more pronounced (as shown in Fig. 4). Notably, the NFW and Beta models exhibit nearly identical orbital trajectories for all dark matter masses $k$ considered in this study, which is consistent with their overlapping in precession angle curves shown in Fig. 5. This demonstrates that despite the different functional forms of NFW and Beta density profiles, they produce nearly indistinguishable effects on both the energetics and shapes of periodic orbits. Furthermore, Fig. 6 demonstrates that periodic orbits calculated in NFW and Beta models are relatively more extended than those in Moore model and Schwarzschild black hole cases. This is caused by the fact that the orbital energies calculated in the NFW and Beta models are substantially lower than those in the Moore model for the same precession angle q (as illustrated in Figs. 4 and 5), since orbital energy is inversely related to the apoapsis distance. This finding suggests a stronger gravitational influence from the NFW/Beta dark matter halos, which is consistent with the results for the ISCO and MBO in 2.2.
@@
Furthermore, to isolate the influences of dark matter halo models, we construct periodic orbits within the NFW model using different values of $k$, each combined with five distinct orbital configurations (zwv)(z~w~v). The results, presented in Fig. 7, clearly demonstrate the effect of dark matter mass k on periodic orbital shapes. For example, the relationship between apoapsis distance and dark matter mass exhibits a consistent monotonic trend: an increase in dark matter mass leads to further apoapsis distance and more pronounced differences of orbital shape, compared with Schwarzschild case. Additionally, from the comparative analysis in Figs. 6 and 7, we observe that the orbital configurations (zwv)(z~w~v) directly determine the geometric shape of periodic orbits. Specifically, (zwv)(z~w~v) represent the number of leaves, the count of additional whirls, and the behavior of the subsequent vertex reached by the particle after leaving the initial vertex (apoapsis). As a result, larger parameter values correspond to more complex orbital structures, which is consistent with the zoom-whirl-vertex classification of periodic orbits in pioneering studies [38, 39].
3.2 The Effect of Dark Matter Halo Scale on Periodic Orbits
@@
4 Gravitational waveforms from periodic orbits
Extreme Mass Ratio Inspiral (EMRI) systems are key target sources for future space-based gravitational wave detectors. These systems are formed by stellar-mass compact objects orbiting around supermassive black holes, and their gravitational wave signals contain rich information about the system’s dynamics and the spacetime geometry of the central black hole. When a small object orbits around a supermassive black hole enveloped by dark matter (DM) in a periodic motion, the gravitational waves it emits provide a unique avenue for studying the system’s properties. This section provides an analysis of the gravitational waveforms of EMRI systems generated by periodic motions, and the theoretical framework for calculating gravitational waves from such periodic orbits is also briefly reviewed. In the calculation of gravitational waves in EMRI systems, the adiabatic approximation and Numerical Kludge waveform model, which is applicable when the small object moves in a nearly static gravitational background with circular, precessional or periodic motion, is adopted [113, 114, 115, 116, 117, 118, 119, 120, 121, 122]. Under this approximation, the energy and angular momentum of the smaller object can be considered constant over several orbital periods, and its trajectory can be viewed as a geodesic in the static background spacetime, with the gravitational radiation’s reaction on the object’s motion temporarily ignored.
The Numerical Kludge waveform model provides a practical scheme for calculating gravitational waves from periodic orbits in a DM halo environment [123]. This method consists of two steps: first, the motion equations Eqs. (2.10a) , (2.10b) and (2.12) are numerically solved to determine the orbit of the small body in the gravitational spacetime containing the DM distribution; then, the quadrupole formula for gravitational radiation is applied to this orbit to generate the corresponding waveform [124]. This method can conveniently reveal the gravitational wave signals of a slowly evolving EMRI system, providing possibilities for exploring the characteristics of the periodic orbits (or precessional orbits), central black hole, and surrounding DM distribution. For the spacetime metric perturbation $h_{ij}$, the gravitational radiation quadrupole moment formula calculated to second order can be expressed as [125, 126, 48],
$h_{ij}=\frac{4\mu M}{D_{L}}\left(v_{i}v_{j}-\frac{m}{r}n_{i}n_{j}\right),$ (4.1)
@@ 4 Gravitational waveforms from periodic orbits
$\mathbf{e}_{X}$ $=(\cos\zeta,-\sin\zeta,0),$ (4.2a)
$\mathbf{e}_{Y}$ $=(\sin\iota\sin\zeta,\cos\iota\cos\zeta,-\sin\iota),$ (4.2b)
$\mathbf{e}_{Z}$ $=(\sin\iota\sin\zeta,-\sin\iota\cos\zeta,\cos\iota).$ (4.2c)
In general relativity, the polarization state of gravitational waves is typically decomposed into two independent modes: ++ (orthogonal) polarization and $\times$ (cross) polarization. In the introduced coordinate system, these two polarization components $h_{+}$ and $h_{\times}$ detected by observers can be expressed as [48, 65],
$h_{+}$ $=\frac{1}{2}(e^{i}_{X}e^{j}_{X}-e^{i}_{Y}e^{j}_{Y})\times h_{ij}=-\frac{2\mu M^{2}}{D_{L}r}\left(1+\cos^{2}\iota\right)\cos\left(2\phi+2\zeta\right)$ (4.3a)
$h_{\times}$ $=\frac{1}{2}(e^{i}_{X}e^{j}_{Y}+e^{i}_{Y}e^{j}_{X})\times h_{ij}=-\frac{4\mu M^{2}}{D_{L}r}\cos\iota\sin\left(2\phi+2\zeta\right),$ (4.3b)
@@
4.1 The effect of dark matter mass on gravitational waves
@@
4.2 The effect of dark matter halo scale on gravitational waves
@@ 5 Conclusion
This study systematically investigates the influence of dark matter halos on periodic orbits around a black hole (with zoom-whirl-vertex behavior) and the produced gravitational wave radiation. This work is carried out within the framework of spherically symmetric dark matter halo models, which have been widely used in astrophysical studies. Through detailed analysis of three dark matter halo models—NFW, Beta, and Moore—we examine how dark matter halo parameters affect spacetime metrics, effective potentials, and properties of periodic orbits around black holes, as well as the resulting gravitational wave signals.
@@
The effects of dark matter halos on spacetime geometry, periodic orbits and gravitational waves can be incorporated into two parameters: the dark matter mass $k$ and halo characteristic radius $h$. Our analysis of the effective potential reveals that the existence of dark matter halos reduces the extrema of the effective potential. This reduction becomes more obvious as the dark matter mass $k$ increases, directly affecting the MBO and ISCO of test particles. The presence of dark matter halos increases the angular momentum for both MBO and ISCO. Particularly, the NFW and Beta models showing nearly identical behavior on ISCO (or MBO), while the Moore model exhibits lower angular momentum values of ISCO (or MBO).
@@
To comprehensively explore periodic orbits, we examine the effects of both dark matter mass $k$ and halo characteristic radius $h$ under the condition of selected angular momentum $L=L_{\text{ISCO}}+\varepsilon(L_{\text{MBO}}-L_{\text{ISCO}})$ and the selected energy $E$ (in Appendix B). For the selected angular momentum condition, it is demonstrated that for a given precession angle $q$, larger dark matter masses require lower orbital energies $E$. The NFW and Beta models require substantially lower energies $E_{(z~w~v)}$ than the Moore model for the same precession angle and orbital configurations (zwv)(z~w~v). This energy difference directly affects the orbital shapes, producing the stretching effects on periodic orbital shapes. An increasing of dark matter mass $k$ greatly stretches the apoapsis of periodic orbits, which leads to larger orbital dimensions. On the other hand, larger dark matter halo scales require higher orbital energy for a same precession angle $q$. At smaller halo scales ($h\sim 10^{7}M$), periodic orbital shapes in dark matter environments significantly differ from the Schwarzschild case. As the halo scale increases, the orbits gradually converge towards the Schwarzschild’s orbital trajectories, closely approaching them at large halo scales. This behavior reflects the dilution of gravitational effects: when the dark matter mass remains constant while the halo scale increases, the gravitational binding becomes weakens in the spacetime. Objects must then possess higher energies (closer to the asymptotic value $\lim_{r\to\infty}V_{\text{eff}}=1$ without gravitation) to maintain the same precession characteristics. This scale-dependent effect is clearly shown and explained with five different periodic orbit configurations (zwv)(z~w~v), where orbits gradually approach the Schwarzschild case as $h$ increases. In particular, when the dark matter halo scale is $h\sim 10^{10}M$, the periodic orbits of all three dark matter models become indistinguishable from the Schwarzschild reference, exhibiting the dilution limit of dark matter halo effects. Furthermore, the analysis of periodic orbits with fixed orbital energy $E$ (in appendix) yields consistent conclusions with those obtained for fixed angular momentum $L$.
@@
To explore the observational signatures of these periodic orbits, we analyze the extreme-mass-ratio inspiral (EMRI) system with a typical mass ratio and luminosity distance. The gravitational wave analysis shows strong correspondence between the gravitational waveforms and orbital properties. Larger dark matter mass leads to more pronounced differences in generated gravitational waves between periodic orbits around SMBH in dark matter halos and those of Schwarzschild black holes, enhancing the distinguishability among different dark matter halo models. The waveforms exhibit stretching trends, particularly an increase in the time duration of each complete orbital period, which stems directly from the modification of the gravitational potential by the dark matter halo. In contrast to the dark matter mass effects, a larger dark matter halo scale $h$ results in a more diffuse mass distribution, weakening its effect on the gravitational potential. The gravitational waveforms generated by periodic orbits exhibit significant differences between various dark matter halo models under a relatively small halo scale ($h\sim 10^{7}M$), which can be used to constrain dark matter models with future observations. The gravitational waveforms gradually approach those for Schwarzschild black holes when dark matter halo scales become sufficiently large. Particularly, when the characteristic radius of the dark matter halo is $h\sim 10^{10}M$, the periodic orbits and their gravitational waveforms converge precisely with those of the Schwarzschild black hole. Furthermore, examining the waveform structure reveals that the orbital configurations (zwv)(z~w~v), as fundamental quantities characterizing the geometric shapes of periodic motion, directly determine the spatiotemporal features (the zoom-whirl stages) of the gravitational wave signals.
@@
Throughout our analysis, we observe that the NFW and Beta models produce nearly indistinguishable results for both orbital characteristics and gravitational wave signals for all parameter ranges examined in this study. This is in accordance with the conclusions of precession angles, effective potentials, and ISCO/MBO, suggesting that these two density profiles have very similar effects on particle orbits and gravitational radiation in the parameter ranges relevant to galactic dark matter halos. The Moore model, while following similar qualitative trends, generates different quantitative results compared with Beta and NFW models, which become more pronounced for larger dark matter masses.
@@
In summary, this study provides an systematic analysis on how dark matter halo parameters influence periodic orbits around black holes and their gravitational wave radiation, through the comparisons of three dark matter halo models. The complementary effects of the dark matter mass and its characteristic radius offer a pathway to understand the complex interplay between dark matter halo and SMBH in galactic center, as well as provide theoretical predictions for gravitational waveforms that can be compared with observations. These findings may provide useful insights and guidance for interpreting gravitational wave signals from EMRIs detected by future space-based observatories, such as LISA, TianQin, and Taiji.